OpenStudy (anonymous):

This is a really easy question, please! The first one to explain it to me and give me the right answer gets a medal. PLEASE HELP SOON!! ~ ~ Samuel is researching the favorite foods of the students in his grade. He surveys the students in his science class. Which of the following is not true?

OpenStudy (anonymous):

Here is the question.

OpenStudy (anonymous):

Did it specify he used stratified sampling?

OpenStudy (anonymous):

He didn't use stratified sampling. To use a stratified sample, you would take, for example, one student from each class in the grade and give them the survey.

OpenStudy (anonymous):

A significant fact.

OpenStudy (anonymous):

lol thanks I got it... but now I have a new question ad;lfjadslkjfa;sdlfj @douglaswinslowcooper @applemuffin95

OpenStudy (anonymous):

Where's your Q?

OpenStudy (anonymous):

Here. @douglaswinslowcooper

OpenStudy (anonymous):

@douglaswinslowcooper The one that is blue and bubbled in is the one I think. I think that's a spinner because their spinners have like, 10 things on it. But a deck of cards have 4 suits. A die has 6 sides. Slips of paper? lol

OpenStudy (anonymous):

You want simulations that easily work with probability of 1 chance in 4. For example, the deck of cards could have each suit correspond to a gumball flavor. Two of the other three could be made to work. the third would be awkward. Do you see which one?

OpenStudy (anonymous):

The spinner? @douglaswinslowcooper

OpenStudy (anonymous):

Spinner could have 4, 8, 12, 16 locations, and so might work. Lots of slips of paper could be used, with a fourth of them the target flavor. How would a die or dice work?

OpenStudy (anonymous):

Spinner, perhaps. Die, certainly.

OpenStudy (anonymous):

OOOhhhh so the answer is a die... @douglaswinslowcooper lol now there is a graph I need help with omg seriously I did all the essays and stuff on my own, and it's 11:16. Can you help too?

OpenStudy (anonymous):

I'm hooked. I'll try.

OpenStudy (anonymous):

haha omg thank you so much. here it is, I'm working on it. you can help too. @douglaswinslowcooper

OpenStudy (anonymous):

@douglaswinslowcooper

OpenStudy (anonymous):

lol I was good with everything else in the module. The only thing that I couldn't do was probability... This part of the test is about probability. ( / )_( \ ) @douglaswinslowcooper

OpenStudy (anonymous):

lol there is another matching thing after this... I think only 1 though. :) I answered like, 3 questions. whoop whoop @douglaswinslowcooper

OpenStudy (anonymous):

noooo omg where did you goooooo I need your help like, really bad. O.O @douglaswinslowcooper

OpenStudy (anonymous):

Fine. We have D M F S, as Darnell, Mother, Father, Sister Without replacement means there are three choices left after the first pick. With replacement means there are four choices again. Note that "parent" means either M or F, an added complication. P(D and S), with replacement First pick: could get D or S out of the group of four, so p1=2/4 Second pick, must get the other one (D if have S. S if have D), so p2=1/4 Joint probability: p1*p2 = 2/16 = 1/8 = .125 My problem is that none of your choices include 1/8, so I'm off or they are. I'll try another.

OpenStudy (anonymous):

if it is soooo simple....

OpenStudy (anonymous):

P(M and F, without replacement): First pick, p1=2/4 as either M or F ok Second pick, p2=1/3, as three are left, one either M or F not gotten before. Joint probability = p1*p2 = 1/12 , one of your choices

OpenStudy (anonymous):

P(parent and sister, without replacement 1st pick M or F: p1 = 1/2 2nd pick, have one parent, D, S left, thus p2 = 1/3 Joint p1*p2= 1/6

OpenStudy (anonymous):

P(parent and child, with replacement): 1st pick M or F out of four: p1 = 2/4 2nd pick D or S out of four: p2 = 2/4 joint = p1*p2 = 4/16 = 1/4

OpenStudy (anonymous):

omg thank you so much, you don't know how much this means. @douglaswinslowcooper

OpenStudy (anonymous):

okay, ummm... I think I only have one more question... lemme check. @douglaswinslowcooper

OpenStudy (anonymous):

uhh I have one more graph... that's it I think. @douglaswinslowcooper

OpenStudy (anonymous):

are you up to it? @douglaswinslowcooper